peanutes worth $2.90 a pound were mixed with cashews worth $4.60 a pound to produce a mixture worth $3.50 a pound. How many pounds of each kind were used to produce 51 pounds of the mixture?
Answer by mananth(16946)

Price ---------------- quantity
Peanuts 2.90 ---------------- x pounds
cashews 4.60 ------ 51 - x pounds
Mixture 3.50 ---------------- 51
Total 51 pounds
2.9 x + 4.6 ( 51 - x ) = 178.5
2.9 x + 234.6 - 4.6 x = 178.5
2.9 x - 4.6 x = 178.5 - -234.6
-1.7 x = -56.1
/ -1.7
x = 33 pounds Peanuts
18 pounds cashews

m.ananth@hotmail.ca